what is the hourly wage for 34k/year?
saxon algebra 2--pg.122

Answer by nerdybill(7011)

what is the hourly wage for 34k/year?
.
34000 1 year 1 week
------ x -------- x -------
year 52 weeks 40 hours
.
34000/(52*40) = 34000/2080 = $16.35/hour